OpenStudy (anonymous):

I WILL GIVE MEDALS!!!! What is the x-intercept of the line with the equation ? 1/3x+y=-15

OpenStudy (dtan5457):

Ok, isolate y. y=-15-1/3x To get the x intercept, y must be 0 0=-15-1/3x solve for x. 15=-1/3x x=?

OpenStudy (anonymous):

how do I do that?? I'm clueless.

OpenStudy (dtan5457):

15=-1/3x what number times -1/3=15..?

OpenStudy (anonymous):

do I turn the fraction into a decimal??

OpenStudy (dtan5457):

you really don't have to. 15=-1/3x divide both sides by -1/3

OpenStudy (anonymous):

5?

OpenStudy (dtan5457):

no, try again...the number should be bigger than 15 AND negative.

OpenStudy (texaschic101):

to divide fractions, flip what you are dividing by, then multiply

OpenStudy (anonymous):

I'm still not getting it. sorry I'm terrible at math.

OpenStudy (texaschic101):

15 = -1/3x --- we have to divide both sides by -1/3 to get x by itself 15/(-1/3) = x -- now flip what you are dividing by and multiply 15 * -3/1 = your answer...can you multiply them ?